OpenStudy (trixitang):

Find the diameter of a cone that has a volume of 29.31 cubic inches and a height of 7 inches. Use 3.14 for pi. (1 point) 4 inches 2 inches 8 inches 6 inches

OpenStudy (trixitang):

@sammixboo help ?

OpenStudy (trixitang):

@mathmate

OpenStudy (sylph):

\[v=πr^{2}\frac{ h }{ 3 }\]

OpenStudy (sylph):

Plug in the volume and the height.

OpenStudy (sylph):

This will go as fast as you do :)

OpenStudy (sylph):

Try to plug them into the formula.

OpenStudy (trixitang):

i have no idea tbh

OpenStudy (sylph):

Put the volume of the cone where the v is and the height where the h is.

OpenStudy (mjdennis):

If you are doing exams by getting help on Open Study, you are violating the OpenStudy Anti-Cheating Policy: http://openstudy.com/terms-and-conditions

OpenStudy (trixitang):

fyi thats what people do here sooo like idc rn

OpenStudy (mjdennis):

@jigglypuff314 As a mod, could you weigh in here? I doubt @trixitang is right, that "what people do here" is cheat on exams.

OpenStudy (sylph):

Oh, plenty of people do. Around this time, most of them are getting answers for semester exams. That's part of the reason why we don't give direct answers: if they're going to cheat, at least make them learn.

OpenStudy (legomyego180):

haha youre such a trixi

jabez177 (jabez177):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Sylph Oh, plenty of people do. Around this time, most of them are getting answers for semester exams. That's part of the reason why we don't give direct answers: if they're going to cheat, at least make them learn. \(\color{#0cbb34}{\text{End of Quote}}\) Hmm. Pretty smooth haha

OpenStudy (legomyego180):

https://www.youtube.com/watch?v=HBFEvDIZpiY

OpenStudy (legomyego180):

@jabez177 how do you quote?

OpenStudy (sylph):

She just used LaTeX to create colored text.

OpenStudy (legomyego180):

ah got it

jabez177 (jabez177):

I used @Jaynator495's extension actually.

jabez177 (jabez177):

It gives us an option. :) http://prntscr.com/bey2jr

OpenStudy (sylph):

That's awesome :3 I just figured the text could be created using latex

OpenStudy (legomyego180):

\(\color{blue}{\text{Originally Posted by}}\) @jabez177 It gives us an option. :) http://prntscr.com/bey2jr \(\color{blue}{\text{End of Quote}}\) very cool!

OpenStudy (trixitang):

wait im so confused, do i have to round up my answer ?

pooja195 (pooja195):

\[v=πr^{2}\frac{ h }{ 3 }\] We are given this much info : \[29.31=(3.14)r^{2}\frac{ 7 }{ 3 }\]

pooja195 (pooja195):

All we need to do is isolate the r^2 any ideas on how we should do that? @trixitang

OpenStudy (trixitang):

okay so the radius is half of the diameter correct ? So the radius is 14.65 correct ? so wouldnt i just multiply that by itself ?

pooja195 (pooja195):

How did you get that?

OpenStudy (trixitang):

how did I get what ?

pooja195 (pooja195):

14.65

OpenStudy (trixitang):

well from dividing the diameter by 2

OpenStudy (sylph):

That's the volume, not the diameter.

OpenStudy (trixitang):

oh then i confused myself horribly

OpenStudy (sylph):

Poo is trying to help you make sense of it. Go along with exactly what she does, and you'll start to understand. Stop doubting yourself and assuming you won't understand.

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 \[v=πr^{2}\frac{ h }{ 3 }\] We are given this much info : \[29.31=(3.14)r^{2}\frac{ 7 }{ 3 }\] \(\color{#0cbb34}{\text{End of Quote}}\) Start by simplfying the things that can be simplified. \[\huge~\rm~3.14 \times \frac{ 7 }{ 3 }=?\]

OpenStudy (trixitang):

7.32 ?

pooja195 (pooja195):

Good let's keep the full decimal though. Now we have \[\huge~\rm~29.31=7.326667r^2\] Any ideas on what to do next?

OpenStudy (trixitang):

I do not

pooja195 (pooja195):

To get a positive outcome Subtract 7.326667r^2 from both sides. 29.31−7.326667r^2=7.326667r^2−7.326667r^2 −7.326667r2+29.31=0 And then you would Subtract 29.31 from both sides. −7.326667r^2+29.31−29.31=0−29.31 \[\huge~\rm~−\frac{ −7.326667r^2=−29.31 }{ −7.326667 }\]From here you divide

pooja195 (pooja195):

Divide both sides by −7.326667

pooja195 (pooja195):

-29.31/−7.326667=?

pooja195 (pooja195):

@trixitang

OpenStudy (trixitang):

4 ?

pooja195 (pooja195):

The answer to the question? (Your question)

OpenStudy (trixitang):

so A ?

pooja195 (pooja195):

yes

OpenStudy (trixitang):

thank you ! Can you help me with one more ?

pooja195 (pooja195):

ill try :)

OpenStudy (trixitang):

okay ill open up a new thread